# The Dark Side of Rich Rewards: Understanding and Mitigating Noise in VLM Rewards

### v4

@@ Abstract
While Vision-Language Models (VLMs) are increasingly used to generate reward signals for training embodied agents to follow instructions, our research reveals that agents guided by VLM rewards often underperform compared to those employing only intrinsic (exploration-driven) rewards, contradicting expectations set by recent work. We hypothesize that false positive rewards – instances where unintended trajectories are incorrectly rewarded – are more detrimental than false negatives. We confirmed this hypothesis, revealing that the widely used cosine similarity metric is prone to false positive estimates. To address this, we introduce BiMI (Binary Mutual Information), a novel reward function designed to mitigate noise. BiMI significantly enhances learning efficiency across diverse and challenging embodied navigation environments. Our findings offer a nuanced understanding of how different types of reward noise impact agent learning and highlight the importance of addressing multimodal reward signal noise when training embodied agents [^1] Code available at https://shorturl.at/VsH70.
The Dark Side of Rich Rewards:
Understanding and Mitigating Noise in VLM Rewards
Sukai Huang1, Shu-Wei Liu2, Nir Lipovetzky1 and Trevor Cohn1thanks: Now at Google DeepMind 1School of Computing and Information Systems, The University of Melbourne, Australia 2 Max Planck Institute for the Physics of Complex Systems, Nöthnitzer Strasse 38, 01187 Dresden, Germany sukaih@student.unimelb.edu.au, sliu@pks.mpg.de, {nir.lipovetzky, trevor.cohn}@unimelb.edu.au
@@ 1 Introduction
Natural language instructions are increasingly recognized as a valuable source of reward signals for guiding embodied agents to learn complex tasks. In particular, a growing trend in embodied agent learning involves using vision-language models (VLMs) for reward modeling. This approach measures the semantic similarity – often quantified by cosine similarity – between the embedding representations of an agent’s behaviors (i.e., past trajectories) and the provided instructions, all within the same embedding space (Kaplan et al., 2017; Goyal et al., 2019, 2020; Du et al., 2023).
Figure 1: Schematic diagram of false positives in a VLM embedding space. The unintended agent’s trajectory (dashed line) may exhibit high cosine similarity to the instruction (solid line) in the embedding space, as indicated by the proximity of their endpoints in the embedding space. Despite this apparent similarity, the unintended trajectory fails to fulfill the instruction. This illustrates how approximation errors can lead to false positive rewards. Refer to Section 5 for more details.
@@
However, we observed that embodied reinforcement learning (RL) agents trained with VLM rewards, while effective in simplified settings, often struggled with tasks involving complex dynamics and longer action horizons. This is evident in several recent works – for instance, Goyal et al. (2019) reported the effective use of VLM rewards in Montezuma’s Revenge, a notoriously challenging Atari game. However, we observed that this success was confined to individual sub-tasks and the agent struggled when attempting to scale up to the full game. Similarly, Du et al. (2023) demonstrated impressive performance of VLM rewards in guiding agents within a 2D survival game. However, their study was conducted in a modified environment with a reduced observation and action space using internal game state information and manually defined macro actions. Consequently, when tested in the original, unmodified environment, their agent’s performance did not exceed that of agents using only intrinsic (exploration-driven) rewards.
@@
The consistent underperformance of VLM rewards, particularly their unexpected failure to outperform intrinsic rewards, raised concerns about their reliability. This discrepancy, where VLM rewards underperformed contrary to their perceived potential, prompted us to investigate the underlying causes of this performance gap. Our findings indicate that noisy reward estimates in VLMs are a key factor contributing to poor learning efficacy. Specifically, our analysis centered around two classes of noise: false positives, which involve rewarding unintended trajectories, and false negatives, which occur when correct trajectories are not rewarded. We posit that false positive rewards (see Figure 1) are not only more prevalent but potentially more detrimental to the learning process than false negatives, a hypothesis supported by our empirical and theoretical findings. Among various sources of reward noise, our study particularly investigates the approximation errors within the commonly used cosine similarity metric. We examine how these errors generate false positive rewards, which in turn hinder learning.
@@
To this end, we propose a novel reward function, BiMI (Binary Mutual Information Reward). It uses binary reward signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to noisy signal sources. Our experiments demonstrate that BiMI significantly improves the learning efficacy of instruction following agents trained by VLM rewards across various challenging environments.
@@ 2 Related Work
Using VLMs as Reward Models. VLMs have been pivotal in robotics studies, serving as reward models that guide agents to follow instructions (Wang et al., 2018; Shridhar et al., 2022; Mahmoudieh et al., 2022). While most research primarily focuses on leveraging VLMs to overcome the challenge of manual reward design for complex tasks (Clark, 2016), the impact of reward noise and its implications for policy convergence rates are often overlooked. As mentioned in Section 1, some work sidesteps the noisy reward problem by accessing internal state information from the game engine as well as providing predefined action macros (Du et al., 2023; Wang et al., 2023), thereby preventing the accumulation of noise over longer horizon. However, understanding the impact of reward noise from VLMs is crucial for developing a reliable language interface for embodied learning agents in real-world applications.
Mitigating Reward Model Failures. Research by Ghosal et al. (2022) and Fu et al. (2024) has introduced methods to counteract unreliable rewards from learned VLMs. These strategies involve employing a parallel exploration-based policy alongside the reward-maximizing policy, thereby reducing reliance on potentially misspecified VLM rewards. Our work contributes to this growing body of research by proposing a novel reward function that directly mitigates the impact of false positive rewards from VLM-based models, complementing approaches that use exploration policies to escape local optima. Furthermore, we tested the synergy of combining our reward function with exploration strategies, demonstrating how these approaches can be integrated for further advancements.
@@ 3 Formal Problem Statement
In this work, we focus on a sparse reward setting, where the agent receives a $+1$ reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$. This sparse reward setting motivates the use of expert instructions and VLMs to provide auxiliary reward signals for more effective RL. Specifically, we have a walkthrough $L$ that breaks down a complex task into $n$ expert-defined sub-tasks, each represented by a natural language instruction that is not necessarily atomic and can encompass multiple finer sub-goals ($L=\{l_{1},l_{2},\ldots,l_{n}\}$). By following these sequential instructions, the agent can navigate from the initial state towards the goal states. A dedicated non-Markovian VLM-based reward model $r^{v}(\tau_{t},l_{m(t)})$ is used to assess how well the agent’s trajectory at current time $t$ fulfills an instruction sentence $l_{m(t)}$. Here, $m(t)$ is a pointer at time step $t$ that indicates the current instruction the agent is trying to complete [^2] The use of non-Markovian reward functions in MDP has been well-established, particularly through the work on reward machines (Icarte et al., 2018; Corazza et al., 2022). For a complete evaluation, we also tested Markovian version of VLM reward function (Pixl2R) in our experiments.. The VLM provides auxiliary rewards by evaluating the semantic similarity between $\tau_{t}$ and $l_{m(t)}$, as illustrated in Figure 2 (see Appendix A.1 for details).
Figure 2: Illustration of reward from a VLM
@@ 4 Theoretical Analysis
In this section, we theoretically show how auxiliary rewards influence policy convergence. Initially, we establish that providing auxiliary rewards that reflect progress towards the goal state typically accelerates the convergence process compared to scenarios where only sparse environmental reward $r^{e}$ is used. Then, we show that the presence of false positive rewards has a more detrimental effect on convergence than false negatives.
@@ 4.1 Auxiliary Rewards and Convergence
With sparse rewards, the gradient landscape is nearly flat, making gradient-based updates indistinguishable from a random walk in parameter space (Antognini and Sohl-Dickstein, 2018). In this problem, we are interested in $D:=\|\theta_{goal}-\theta_{0}\|$ which is the distance in parameter space from initial parameters $\theta_{0}$ to goal parameters $\theta_{goal}$. We make the following assumption
@@
We also prove the following proposition
@@
The sum of expected time for a series of random walks, each covering the shorter distance of an individual sub-task, is less than the expected time to travel the entire distance $D$ in one long random walk:
$\frac{1}{n-1}\mathbb{E}[T_{D}]\leq\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]<\mathbb{E}[T_{D}].$
and thus show that subgoal-based auxiliary rewards improve the convergence of random walk optimization in a sparse reward landscape up to a factor of $(n-1)$ with details in Appendix A.2.1.
@@ 4.2 Connection to Heuristic-Guided RL
The problem and the possible solutions can be framed within the context of Heuristic-Guided Reinforcement Learning (HuRL) by Cheng et al. (2021). HuRL mandates that auxiliary reward signals serve as heuristics, where $h:\mathcal{S}\rightarrow\mathbb{R}$ approximates the future total rewards an agent expects to get starting from state $s$ under the optimal policy $\pi^{*}$ (i.e., $h(s)\approx V^{*}(s)$). These heuristics typically come from domain knowledge, aligning with expert instructions. We make the following assumption
@@
The fulfillment of $l_{m(t)}$ can be captured by intermediate goal states $\mathcal{S}^{m(t)}$.
This assumption is justified because often instructions in VLM come with end goals that can be indicated by definite states. For further discussions, please refer to Section 9. This assumption is required so that we can construct a domain-knowledge Markovian heuristic reward based on non-Markovian VLM rewards (see Appendix A.1.3 for the exact formula). With this construction, we are estimating the possible reward in the future by evaluating how well the agent has followed expert instruction in the past. It is consistent with the intuition of a heuristic reward based on semantic similarity because the better the agent has been following expert instructions, the more likely it is on the right track to collect more future rewards.
Using HuRL framework allows us to analyze how false positive rewards influence on policy performance gap, defined as $V^{*}(s_{0})-V^{\pi}(s_{0})$. We will demonstrate that false positive rewards increase the upper bound of this gap, whereas false negative rewards maintain this upper bound.
4.3 False Positives vs. False Negatives
To begin, we provide the formal definition of false positive auxiliary rewards from both instruction-following (IF) and heuristic perspectives:
@@
IF Perspective: For a trajectory $\tau_{t}$ that does not satisfy instruction $l_{m(t)}$, the VLM-based reward $r^{v}(\tau_{t},l_{m(t)})$ ranges between 00 and $1$.
@@
The heuristic is approximated as $h(s_{t})\approx V^{*}(s_{t})=1\cdot\gamma^{\widetilde{T}-t}$, calculated based on an assumed optimal path length $\widetilde{T}$. Here, $\widetilde{T}-t$ measures the remaining steps towards the goal. When a trajectory $\tau_{t}$ receives a high reward but fails to fulfill instruction $l_{m(t)}$, it corresponds to a high $h(s_{t})$, thus a low $\widetilde{T}-t$. However, since the agent must eventually redo $l_{m(t)}$, $s_{t}$ is actually further from the goal than estimated. Therefore, the actual distance $T-t$ to reach the goal will exceed $\widetilde{T}-t$, and $V^{*}(s_{t})$, calculated with the actual $T$, will be smaller than $h(s_{t})$, derived from $\widetilde{T}$. This thus explains how a false positive reward from an instruction-following (or VLM) perspective leads to an overestimation of the heuristic. Researchers have advocated the benefits of pessimistic value estimation to enhance the stability of RL algorithms (Kumar et al., 2020; Jin et al., 2021). In HuRL, Cheng et al. (2021) further identify a beneficial property: when a heuristic is pessimistic with respect to $\mathcal{M}$, it results in a smaller upper bound on the performance gap.
Definition 4.6 (Pessimistic $h$).
Let Bellman operator $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. A heuristic function $h$ is said to be pessimistic with respect to an MDP $\mathcal{M}$ if $\max_{a}(\mathcal{B}h)(s,a)\geq h(s)$. This condition essentially means that the heuristic $h$ never overestimates the true value of a state.
@@
Even if the heuristic remains conservative for all successor states, a single overestimation ($h(s)>V^{*}(s)$) can violate the pessimistic condition by causing $max_{a}(\mathcal{B}h)(s,a)<h(s)$.
Informally, our result on the impact of false positive/negative rewards on convergence is as follows:
@@
Borrowing from Cheng et al. (2021), the performance gap in RL when using heuristics can be broken down into a regret term and a bias term, where false negatives maintain the upper bound of bias while false positives increase the bias by breaking the heuristic’s pessimism, thereby potentially leading to slower convergence.
See Appendix A.2.2 for formal version and detailed proofs. In the next section, we present a case study on cosine similarity metrics, demonstrating how they contribute to false positive rewards in learned reward models.

### v5

@@ Abstract
While Vision-Language Models (VLMs) are increasingly used to generate reward signals for training embodied agents to follow instructions, our research reveals that agents guided by VLM rewards often underperform compared to those employing only intrinsic (exploration-driven) rewards, contradicting expectations set by recent work. We hypothesize that false positive rewards – instances where unintended trajectories are incorrectly rewarded – are more detrimental than false negatives. We confirmed this hypothesis, revealing that the widely used cosine similarity metric is prone to false positive estimates. To address this, we introduce BiMI (Binary Mutual Information), a novel reward function designed to mitigate noise. BiMI significantly enhances learning efficiency across diverse and challenging embodied navigation environments. Our findings offer a nuanced insight of how different types of reward noise impact agent learning and highlight the importance of addressing multimodal reward signal noise when training embodied agents.
@@ 1 Introduction
Natural language instructions are increasingly recognized as a valuable source of reward signals for guiding embodied agents to learn complex tasks. In particular, a growing trend in embodied agent learning involves using vision-language models (VLMs) for reward modeling. This approach measures the semantic similarity – often quantified by cosine similarity – between the embedding representations of an agent’s behaviors (i.e., past trajectories) and the provided instructions, all within the same embedding space (Kaplan et al. 2017; Goyal, Niekum, and Mooney 2019, 2020; Du et al. 2023; Rocamonde et al. 2024; Wang et al. 2024).
Figure 1: Illustration of embodied RL agents using VLM reward model
@@
However, we observed that embodied RL agents trained with VLM rewards, while effective in simplified settings, often struggled with tasks involving complex dynamics and longer action horizons [^1] Code available at https://shorturl.at/3uVGP. This is evident in several recent works – for instance, Goyal, Niekum, and Mooney (2019) reported the effective use of VLM rewards in Montezuma’s Revenge, a notoriously challenging Atari game. However, we observed that this success was confined to individual sub-tasks and the agent struggled when attempting to scale up to the full game. Similarly, Du et al. (2023) demonstrated impressive performance of VLM rewards in guiding agents within a 2D survival game. However, their study was conducted in a modified environment with a reduced observation and action space using internal game state information and manually defined macro actions (see Appendix B). Consequently, when we tested their methods in the original, unmodified environment, we found that their agent’s performance did not exceed that of agents using only intrinsic (exploration-driven) rewards.
The consistent underperformance of VLM rewards, particularly their unexpected failure to outperform intrinsic rewards, raised concerns about their reliability. This discrepancy, where VLM rewards underperformed contrary to their perceived potential, prompted us to investigate the underlying causes of this performance gap.
@@
Our findings indicate that noisy reward estimates in VLMs are a key factor contributing to poor learning efficacy. Specifically, our analysis centered around two classes of noise: false positives, which involve rewarding unintended trajectories and can mislead the agent into reinforcing suboptimal behaviors; and false negatives, which occur when correct trajectories are not rewarded, thus providing little feedback and reintroduce the reward sparsity issue. We posit that false positive rewards are not only more prevalent but potentially more detrimental to the learning process than false negatives, a hypothesis supported by our empirical and theoretical findings. Among various sources of reward noise, our study particularly investigates the approximation errors within the commonly used cosine similarity metric. We examine how these errors generate false positive rewards, which in turn hinder learning.
@@
To this end, we propose a novel reward function, BiMI (Binary Mutual Information Reward). It uses binary signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to noisy signal sources. Our experiments demonstrate that BiMI significantly improves the learning efficacy of instruction-following agents trained with VLM rewards across various complex, long-horizon tasks in embodied environments. We validate the effectiveness of BiMI in both Markovian and non-Markovian reward settings, showcasing its robustness and generalizability across different types of reward models.
@@ 2 Related Work
We provide a brief overview of the related work for this study. For detailed ones, please refer to Appendix B.
VLMs as Reward Models. Despite the rising use of VLM-based reward models in embodied RL (Kaplan et al. 2017; Goyal, Niekum, and Mooney 2019, 2020; Du et al. 2023; Rocamonde et al. 2024; Wang et al. 2024, 2019; Shridhar et al. 2022; Mahmoudieh, Pathak, and Darrell 2022), many works focus on proof-of-concept results in simplified environments, sidestepping the challenges posed by noisy reward signals. Critical issues such as reward noise and its effect on policy convergence remain largely understudied. In particular, some approaches sidestep noise by leveraging internal state information or predefined action macros (Du et al. 2023; Wang et al. 2023). However, these methods struggle to scale to real-world applications due to their reliance on environment-specific priors, which lack generalizability in dynamic, unstructured settings.
Pessimistic Value Estimation. The RL community has demonstrated that pessimism under uncertainty improves robustness in settings like offline RL optimization (Uehara and Sun 2021; Bai et al. 2024). These methods conservatively estimate state values to avoid overoptimism in poorly explored or uncertain regions. However, such techniques are rarely applied to settings with learned VLM-based rewards, where noise arises not from environmental uncertainty but from misaligned or overconfident reward signals. Moreover, these methods remain largely alien to recent studies of VLM reward + RL embodied agent research. Our work bridges this gap by applying pessimism to mitigate learned reward noise, addressing a critical yet overlooked challenge in this emerging field.
Optimistic Exploration in Online RL. Pessimistic estimates in RL are mostly an offline phenomenon; online RL instead typically embraces optimism under uncertainty to drive efficient exploration. Classic model-based schemes such as RMax initialize all unknown state-action pairs with the maximum possible reward, thereby provably encouraging the agent to visit under-explored regions of the MDP (Brafman and Tennenholtz 2002). In the model-free regime, intrinsic reward methods measure the surprise of an agent’s actions, such as the prediction error of a forward model (Burda et al. 2018; Wan et al. 2023), or the novelty of state visits (Zhang et al. 2021). Our work investigates the integration of pessimistic, VLM-based reward signals with optimistic intrinsic rewards in an online embodied environment.
Reward Signal from Human Preference. Recent work on RL from Human Feedback (RLHF) (Ouyang et al. 2022) also leverage expert preference as a reward signal. However, our work differs in key aspects. Unlike RLHF’s focus on textual outputs, our approach involves evaluating cross-modal similarities between visual and textual data in environments requiring long-horizon decision-making and frequent embodied interactions, a domain not typically covered by RLHF.
Mitigating Misspecified Rewards. Prior works proposed mitigating false positive rewards by training a parallel exploration policy to escape local optima caused by misspecified rewards (Ghosal et al. 2022; Fu et al. 2024). In contrast, we propose a novel reward function that directly penalizes likely false positive reward signals during training. We further show that our method complements exploration-based methods and achieves superior performance when combined.
@@ 3 Formal Problem Statement
In this work, we focus on a sparse reward setting, where the agent receives a +1+1 reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$. This sparse reward setting motivates the use of expert instructions and VLMs to provide auxiliary reward signals for more effective RL. Here, it’s crucial to note that by VLM, we specifically mean encoder-based VLMs, as opposed to generative VLMs (see Appendix M). For conciseness, we will simply use ‘VLM’ to refer to encoder-based VLMs.
Specifically, we have a walkthrough $L$ that breaks down a complex task into $n$ expert-defined sub-tasks, each represented by a natural language instruction that is not necessarily atomic and can encompass multiple finer sub-goals ($L=\{l_{1},l_{2},\ldots,l_{n}\}$). By following these sequential instructions, the agent can navigate from the initial state towards the goal states. A dedicated non-Markovian VLM-based reward model [^2] The use of non-Markovian reward functions in MDP has been well-established, particularly through the work on reward machines (Icarte et al. 2018; Corazza, Gavran, and Neider 2022). For a complete evaluation, we also tested Markovian version of VLM reward function (Pixl2R) in our experiments. $r^{v}(\tau_{t},l_{m(t)})$ is used to assess how well the agent’s trajectory at current time $t$ fulfills an instruction sentence $l_{m(t)}$. For the detailed mechanism and the pseudo-code of the existing VLM reward + RL algorithm, please refer to Appendix C. Note that our proposed BiMI reward function uses a different mechanism to determine if instruction $l_{m(t)}$ is completed at time $t$ (explained shortly in Section 7). The VLM provides auxiliary rewards by evaluating the semantic similarity between $\tau_{t}$ and $l_{m(t)}$, as illustrated in Figure 1. The use of VLM-based reward model transform the original MDP into a shaped MDP $\widetilde{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},\tilde{r},\gamma\rangle$, where the reward function becomes $\tilde{r}=r^{e}(s,a)+r^{v}(\tau_{t},l_{m(t)})$.
@@ 4 Theoretical Analysis
In this work, we define “reward noise” as errors made by the VLM-based reward model when evaluating an agent’s trajectory: either spuriously rewarding a trajectory that doesn’t satisfy the instruction (false positive) or failing to reward one that does (false negative). In this section, we first show that, in the absence of noise, auxiliary rewards derived from language instructions, reflecting progress toward the goal, accelerate convergence compared to relying solely on sparse environmental rewards, $r^{e}$. We then prove that false positive rewards impede convergence more severely than false negatives under the Heuristic guided RL (HuRL) framework (Cheng et al. 2021).
@@ 4.1 Auxiliary Rewards and Convergence
With sparse rewards, the gradient landscape is nearly flat, making gradient-based updates indistinguishable from a random walk in parameter space (see Proposition N.1 in Appendix N). In this problem, we are interested in $D:=\|\theta_{goal}-\theta_{0}\|$ which is the distance in parameter space from initial parameters $\theta_{0}$ to goal parameters $\theta_{goal}$. We make the following assumption:
@@
We therefore prove the following proposition:
@@
The sum of expected time for a series of random walks, each covering the shorter distance of an individual sub-task, is less than the expected time to travel the entire distance $D$ in one long random walk: $\frac{1}{n-1}\mathbb{E}[T_{D}]\leq\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]<\mathbb{E}[T_{D}]$.
and thus show that subgoal-based auxiliary rewards improve the convergence of random walk optimization in a sparse reward landscape up to a factor of (n1)(n-1). See Appendix D for the proof. The theoretical takeaway is that auxiliary rewards aligned with intermediate subgoals can accelerate convergence of RL algorithms compared to learning from a single delayed reward at the goal state. In practice, however, this improvement depends critically on the quality of the auxiliary reward signal — which we examine in the next section.
@@ 4.2 Connection to Heuristic-Guided RL
The reward noise issue can be framed within the context of HuRL by Cheng et al. (2021). HuRL mandates that auxiliary reward signals serve as heuristics, where $h:\mathcal{S}\rightarrow\mathbb{R}$ approximates the future total rewards an agent expects to get starting from state $s$ under the optimal policy $\pi^{*}$ (i.e., $h(s)\approx V^{*}(s)$). First of all, we establish the following assumption:
@@
The expert instruction sequence $L={l_{1},l_{2},\dots,l_{n}}$ specifies a sequence of state and action landmarks guiding the agent toward goal states $\mathcal{S}_{G}$. Landmarks are key states or key actions that must hold or occur in any trajectories that reach $\mathcal{S}_{G}$.
Assumption 4.3 is reasonable and reflects the quality of expert instructions, which are crucial for the agent to reach the goal efficiently.
We now present the important proposition connecting VLM-based reward models to HuRL:
Proposition 4.4.
In sparse reward settings, VLM-based reward models, as implemented in Algorithm 1, can be viewed as a heuristic function $h(s_{t})$ in HuRL that estimates $V^{*}(s_{t})$.
The proof is in Appendix E. HuRL framework allows us to analyze how false positive rewards influence on optimality gap, defined as $V^{*}(s_{0})-V^{\pi}(s_{0})$, a metric that is used for theoretical analysis of the convergence speed of RL algorithms — the smaller the gap, the faster the convergence. We demonstrate that false positive rewards increase the upper bound of this gap, whereas false negative rewards maintain this upper bound.
4.3 False Positives and Heuristic Overestimation
We provide the formal definition of false positive and false negative rewards from both instruction-following (IF) and heuristic perspectives:
@@ 4.3 False Positives and Heuristic Overestimation
IF Perspective: For a trajectory $\tau_{t}$ that does not satisfy instruction $l_{m(t)}$, the VLM-based reward $r^{v}(\tau_{t},l_{m(t)})>0$.
@@
Definition 4.6 (False Negative Rewards).
A false negative reward occurs when:
IF Perspective: The VLM-based reward $r^{v}(\tau_{t},l_{m(t)})\approx 0$ for a trajectory $\tau_{t}$ that does satisfy instruction $l_{m(t)}$.
@@
Proof.
The heuristic is approximated as $h(s_{t})\approx V^{*}(s_{t})=1\cdot\gamma^{\widetilde{T}-t}$ in sparse reward setting, calculated based on an assumed optimal path length $\widetilde{T}$. Here, $\widetilde{T}-t$ measures the remaining steps towards the goal. When a trajectory $\tau_{t}$ receives a high reward but fails to fulfill instruction $l_{m(t)}$, it corresponds to a high $h(s_{t})$, thus a low $\widetilde{T}-t$. Yet the agent must eventually reattempt $l_{m(t)}$ due to Assumption 4.3, $s_{t}$ is actually further from the goal than estimated. Therefore, the actual distance TtT-t to reach the goal will exceed $\widetilde{T}-t$, and $V^{*}(s_{t})$, calculated with the actual $T$, will be smaller than $h(s_{t})$, calculated with $\widetilde{T}$. This thus explains how a false positive reward from an instruction-following (or VLM) perspective leads to an overestimation of the heuristic. ∎
Researchers have advocated the benefits of pessimistic value estimation to enhance the stability of RL algorithms (Kumar et al. 2020; Jin, Yang, and Wang 2021). In HuRL, Cheng et al. (2021) further identify a beneficial property: when a heuristic is Bellman-consistent pessimism (i.e., $\max_{a}(\mathcal{B}h)(s,a)\geq h(s)$), it results in a capped upper bound on the optimality gap.
Definition 4.8 (Bellman-consistent Pessimistic $h$).
Recall Bellman equation of $h$: $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$. A heuristic function $h$ is said to be Bellman-consistent pessimism with respect to an MDP $\mathcal{M}$ if $\max_{a}(\mathcal{B}h)(s,a)\geq h(s)$. This condition essentially means that the heuristic $h$ never overestimates the true value of a state.
@@
Even if the heuristic remains conservative for all successor states, a single false positive overestimation, i.e., $h(s)>V^{*}(s)$, can violate the pessimistic condition by causing $max_{a}(\mathcal{B}h)(s,a)<h(s)$.
The proof is in Appendix F. This implies that maintaining a pessimistic heuristic is inherently fragile, as the presence of a false positive in any state disrupts the pessimistic condition.
Our result on the impact of false positive and false negative rewards on convergence is captured as follows:
@@
Adapting the optimality gap analysis from Cheng et al. (2021), the gap $V^{*}(s_{0})-V^{\pi}(s_{0})$ in heuristic-guided RL decomposes into regret and bias terms. False negatives preserve the upper bound on bias, as they maintain the heuristic’s pessimism, i.e., $h(s)\leq V^{*}(s)$. In contrast, false positives breaks the pessimistic property, strictly increase the bias without an upper bound, thereby leading to a larger optimality gap and slower convergence.
See Appendix G for the formal version and detailed proof. In the next section, we present a case study on cosine similarity metrics, demonstrating how they contribute to false positive rewards in learned reward models.